OpenStudy (anonymous):

The total number of ways in which 5 balls of different colours can be distributed among 3 persons so that each persons gets at least one ball is?

OpenStudy (shubhamsrg):

let no. of balls each person gets be p1,p2,p3 then we have p1 + p2 + p3 = 5 where p1 >=1 p2>=1 and p3>=1 okay,,before that,, suppose we had this a + b+c ....r = n where each a,b,c >=0 no. of such arrangements = C(n+r -1 , r-1) it'd be better if you directly rote down this formulla,,otherwise this thing can also be explained..needless for iit preparation though.. now..what we have is p1,p2,p3 >=1 and not 0 so we convert that we can see p1 - 1 >=0 p2- 1 >=0 p3 -1 =0 we had p1+p2+p3 =5 or (p1 -1) + (p2 -1) + (p3 -1) = 2 so no. of ways will be C(2+3-1 , 3-1) =C(4,2) = 6